# A Survey of Low-bit Large Language Models: Basics, Systems, and Algorithms

### v2

@@ 5.2.1 Equivalent Transformation
However, both the orthogonal matrices in QuIP and the Hadamard matrices in QuIP# and QuaRot are randomly generated. Although these works have shown that these randomly generated matrices can alleviate the outlier problem to some extent, they are not optimal. SpinQuant (Liu et al., 2024e) finds that the performance of a quantized network can vary significantly with different rotation matrices. For example, the average accuracy on downstream zero-shot reasoning tasks can fluctuate by up to 13 points depending on the rotation used on the MMLU benchmark. Therefore SpinQuant proposes a learning-based rotation transformation. The rotation matrix is learned using the Cayley SGD method, with the following optimization objective:
@@
Here, $\mathcal{M}$ presents the Stiefel manifold, i.e., the set of all orthogonal matrices. $\mathcal{L}_{Q}(\cdot)$ denotes the task loss. By employing the learned matrix, the performance is improved significantly and the variance becomes much smaller compared with randomized matrices. The diagram in SpinQuant (Liu et al., 2024e) effectively illustrates the overall process of the rotation transformation, so we have borrowed it for our purposes as shown in Figure 11.
We can observe that scaling transformation and rotation transformation can be utilized for the different parts of LLM quantization. QServe (Lin et al., 2024b) is a co-designed quantization system for efficient LLM serving, combining scaling and rotation transformations. For where additional overhead is required for the online computation of rotation matrices, QServe uses scale transformation as a substitute for rotation operations, thereby avoiding the extra overhead.
@@ 5.2.2 Compensation
The weight compensation technique, originally stemming from Optimal Brain Damage (OBD) (LeCun et al., 1989), involves a Taylor series expansion of the objective function. This method assumes that upon the removal of any given parameter, the influence of the remaining parameters on the objective function remains unchanged. Based on OBD, OBS (Hassibi et al., 1993) and OBQ (Frantar and Alistarh, 2022) calculate the impact of each parameter weight on the objective function by solving the inverse Hessian matrix. Concurrently, they compute a compensation term applied to the remaining weights to offset the error introduced by each weight adjustment.
Although one-by-one weight quantization methods have achieved satisfactory performance on smaller models, the computational overhead becomes prohibitive when scaling to larger models. To accelerate quantization, GPTQ (Frantar et al., 2022) quantizes the weights column-by-column, and the rounding errors are compensated using second-order information. Specifically, this algorithm compensates for the quantization error induced by the quantized weights $Quant(\mathbf{W}_{i})$ by adjusting the subset R of full-precision weights $R$ with an update $\boldsymbol{\delta}_{R}$:
@@
where the Hessian matrix is $\mathbf{H}_{R}=2\mathbf{X}_{R}\mathbf{X}_{R}^{\top}$. Based on GPTQ, several works have been successively proposed. QuantEase (Behdin et al., 2023) utilizes the Coordinate Descent to compute more precise compensation for the unquantized weights. QQQ (Zhang et al., 2024e) adopts the GPTQ for the transferred weights by OS+ (Wei et al., 2023b).
@@ 5.2.3 Mixed-precision
As aforementioned, the presence of outliers is widely found in the activations and weights of large language models, which poses a significant challenge for quantization. Consequently, the motivation of numerous mixed-precision methods for LLMs is to represent a small number of outlier values in higher precision and other values in lower precision separately. Similarly, depending on the granularity of mixed precision, methods can be categorized into element-wise, channel-wise, and tensor-wise approaches, as described in Section 2.2.
@@
Element-wise. SpQR (Dettmers et al., 2023) was the first to demonstrate that outliers also exist in weights. It identifies and isolates these outlier weights based on their sensitivity, saving them as a highly sparse, higher-precision matrix. SqueezeLLM (Kim et al., 2023) adopts non-uniform quantization for non-salient weights, which achieves near-lossless performance. Similarly, CherryQ (Cui and Wang, 2024) defines heterogeneity to identify the critical cherry parameters. To explore extreme compression rate, PB-LLM (Shang et al., 2023) is the first to binarize the non-salient weights in LLMs. Since PB-LLM still allocates high precision to 10%-30% of salient weights, BiLLM (Huang et al., 2024) employs residual approximation for salient weights and group quantization for non-salient weights, reducing the quantization bit-width of LLM weights to 1.08 bits. GEAR (Kang et al., 2024) extends the concept of mixed precision to the KV cache compression and utilizes low-rank matrices to approximate the quantization residuals.
@@
Channel-wise. LLM.int8() (Dettmers et al., 2022a) splits the weights and activations into two independent parts according to the outlier channels to minimize output quantization errors in activations, which effectively reduces the GPU memory usage during inference. OWQ (Lee et al., 2024a) proposes a sensitivity-aware mixed-precision scheme to identify the weak columns by Hessian metric. Furthermore, OWQ also provides weak column tuning (WCT) to enable accurate parameter-efficient fine-tuning for task-specific adaptation. RPTQ (Yuan et al., 2023) observes the varying ranges across channels in activations pose challenges for quantization. Therefore, RPTQ reorders the channels into different clusters with respective quantization. Atom (Zhao et al., 2024) employs dynamic reorder for activations and static reorder for weights to remain aligned with the corresponding activation channels. Atom further quantizes the KV cache to 4-bit which significantly boosts serving throughput. Inspired by information theory, CQ (Zhang et al., 2024b) couples multiple key/value channels together and jointly quantize them.
Tensor-wise. DecoQuant (Liu et al., 2024b) aims to migrate the quantization difficulties by performing the tensor decomposition to local tensors. The benchmark for MOE PTQ (Li et al., 2024b) investigates the weight bits among different blocks, experts, and linear layers, which reveals the varying numbers of weight bits are effective.
@@ 5.2.4 Combination
Although QAT is generally considered to offer the best accuracy, its high memory cost makes it difficult to apply to LLMs. Therefore, some methods consider introducing LoRA or other matrix decomposition methods as a trade-off between PTQ and QAT. Unlike PEFT discussed in Section 3.3, these methods aim to reduce quantization error using techniques like LoRA or SVD to achieve a quantized model closer to the full-precision model, rather than enhancing learning ability on fine-tuning datasets. Some works have used LoRA to achieve parameter-efficient QAT. LR-QAT (Bondarenko et al., 2024) computes $s\cdot\mathrm{clamp}(\textbf{W}_{q}+\textbf{A}^{i\times r}\textbf{B}^{r\times o})$ during the forward pass and does not update W during the backward pass, allowing a 7B LLM to be trained on a single consumer-grade GPU with 24GB of memory. This approach results in a quantization-friendly model after fine-tuning. LLM-QFA aims to produce models with various bit widths through a single supernet training, significantly reducing the resource overhead of this production method by leveraging the low resource cost of LoRA. INT2.1 (Chai et al., 2023) utilizes LoRA to shift the optimization target from minimizing per-layer or per-block quantization error to minimizing the overall output error of the model. Through end-to-end fine-tuning, it reduces the distance between the output distribution and its corresponding original full-precision counterpart. Other works have reduced quantization errors through matrix decomposition. LQER (Zhang et al., 2024a) applies SVD to quantization errors and uses an activation-induced scaling matrix to guide the singular value distribution toward the desired pattern. Delta-CoMe (Ping et al., 2024) discovers that the singular values of delta weights exhibit a long-tailed distribution after applying SVD, and proposes a mixed-precision delta quantization method that uses high-bit representations for the singular vectors corresponding to these singular values. ZeroQuant-V2 (Yao et al., 2023) introduced an optimized low-rank compensation method that enhances model quality recovery by leveraging a low-rank matrix obtained through SVD of the quantization errors. LCQ (Cai and Li, 2024) uses low-rank codebooks with a rank greater than one for quantization, addressing the issue of accuracy loss when using rank-one codebooks under high compression ratios.
@@
Model sparsification aims to remove unimportant weights to accelerate the model, while quantization further reduces the remaining weights using lower-bit representations. Therefore, the two methods can be effectively used in a complementary manner. SDQ (Jeong et al., 2024) first sparsifies the weights of LLMs based on the magnitude as much as possible until the quality of the LLM is significantly impacted (e.g., a 1% increase in perplexity). Then it utilizes a mixed-precision quantization method to deal with the outliers. However, this method does not take into account the coupling of the two approaches. Sparsification and quantization often conflict with each other. Sparsification tends to preserve parameters with large absolute values in LLMs (Han et al., 2015; Sun et al., 2023), while quantization prefers a smaller range of parameter values (Wei et al., 2023b). As a result, the parameters preserved during sparsification may degrade the performance of quantization. JSQ (Guo et al., 2024) design a new sparsity metric to address this issue:
$\textbf{I}_{ij}$ $=||\textbf{X}||_{2}\cdot||\textbf{W}||,$ (34)
@@
$where\quad\hat{\textbf{Y}}=\textbf{X}\cdot(\Theta(\textbf{W};i;j))^{\mathsf{T}},$
@@
In addition to combining quantization with other compression methods, different quantization techniques can also be integrated to achieve better results. A recent work (Sharify et al., 2024) combines the SmoothQuant and GPTQ together. Actually, most of the equivalent transformation methods and the compensation quantization methods are orthogonal which can be merged for further exploration.
@@ 5.2.5 More Quantization Forms
Integer quantization typically assigns a single scaling factor to an entire block and quantizes each element individually into an integer number. This reduces memory usage while also enabling the acceleration of fixed-point operations after quantizing both weights and activations. However, as higher precision is demanded for LLM quantization, formats that better match the original distribution of values have been proposed. Normal Float (Dettmers et al., 2021, 2024), proposed alongside Quantile Quantization, is based on the assumption that the weight distribution follows a normal distribution. It is considered an information-theoretically optimal data type that ensures each quantization bin has an equal number of values assigned from the input tensor. However, Dotzel et al. (Dotzel et al., 2024b) conducted a statistical analysis and found that the distributions of most LLM weights and activations follow a Student’s t-distribution. Based on this, they derived a new theoretically optimal format, Student Float (SF4). Floating-Point (FP) quantization offers better hardware support compared to NF/SF and is more flexible than integer quantization, allowing it to more effectively handle long-tail or bell-shaped distributions. Since FP can support flexible allocation of exponent and mantissa bits, several allocation schemes have been proposed. FPQ (Liu et al., 2023a) determines FP quantizers through a joint format and max value search combined with a pre-shifted exponent bias. FP8 quantization (Kuzmin et al., 2022) tests various allocation schemes by evaluating metrics like quantization error and proposes FP8 quantization simulation for learnable allocation and quantization.
@@
Transformer-VQ (Lingle, 2023) applies vector quantization (VQ) to the key vector sequence of Attention, reducing the complexity of Attention to linear. Most other VQ works focus on optimizing the codebooks $C_{m}\in\mathbb{R}^{2^{B}}$, and the discrete codes represented by one-hot $d$. AQLM (Egiazarian et al., 2024b) learns additive quantization of weight matrices in an input-adaptive fashion and jointly optimizes codebook parameters across each transformer block. QuIP# (Egiazarian et al., 2024b) uses vector quantization to exploit the spherical sub-Gaussian distribution inherent in incoherent weights by introducing a hardware-efficient codebook based on the highly symmetrical E8 lattice. GPTVQ (van Baalen et al., 2024) interleaves the quantization of one or more columns with updates to the remaining unquantized weights, using information from the Hessian of the per-layer output reconstruction MSE, and further compresses the codebooks by using integer quantization and SVD-based compression. PV-Tuning (Malinovskii et al., 2024) notes that using straight-through estimators (STE) leads to suboptimal results and proposes an alternating iterative optimization strategy for scales, codebooks, zeros (continuous parameters), and assignments (discrete codes) during fine-tuning. QTIP (Tseng et al., 2024b) uses a stateful decoder that separates the codebook size from the bitrate and effective dimension to achieve ultra-high-dimensional quantization.

### v3

@@ 5.2.1 Equivalent Transformation
However, both the orthogonal matrices in QuIP and the Hadamard matrices in QuIP# and QuaRot are randomly generated. Although these works have shown that these randomly generated matrices can alleviate the outlier problem to some extent, they are not optimal. SpinQuant (liu2024spinquant) finds that the performance of a quantized network can vary significantly with different rotation matrices. For example, the average accuracy on downstream zero-shot reasoning tasks can fluctuate by up to 13 points depending on the rotation used on the MMLU benchmark. Therefore SpinQuant proposes a learning-based rotation transformation. The rotation matrix is learned using the Cayley SGD method, with the following optimization objective:
@@
Here, $\mathcal{M}$ presents the Stiefel manifold, i.e., the set of all orthogonal matrices. $\mathcal{L}_{Q}(\cdot)$ denotes the task loss. By employing the learned matrix, the performance is improved significantly and the variance becomes much smaller compared with randomized matrices. The diagram in SpinQuant (liu2024spinquant) effectively illustrates the overall process of the rotation transformation, so we have borrowed it for our purposes as shown in Figure 11. Specifically, for Quarot (ashkboos2024quarot), since it employs a head-wise rotation transformation at $R_{2}$, an online Hadamard matrix needs to be inserted before quantizing the attention output to achieve an equivalent transformation. DuQuant (lin2024rotation) identifies the limitations of these methods in smoothing massive outliers and therefore utilizes rotation and permutation transformations based on prior knowledge. Meanwhile, unlike SpinQuant, a greedy search strategy is employed to optimize the rotation matrix. PrefixQuant (chen2024prefixquant) discovers the token-wise outliers, especially appearing in initial tokens and low-semantic tokens. Since these tokens remain unchanged across all inputs, PrefixQuant stores their KV cache through offline prefilling.
We can observe that scaling transformation and rotation transformation can be utilized for the different parts of LLM quantization. QServe (lin2024qserve) is a co-designed quantization system for efficient LLM serving, combining scaling and rotation transformations. For where additional overhead is required for the online computation of rotation matrices, QServe uses scale transformation as a substitute for rotation operations, thereby avoiding the extra overhead.
@@ 5.2.2 Compensation
The weight compensation technique, originally stemming from Optimal Brain Damage (OBD) (lecun1989optimal), involves a Taylor series expansion of the objective function. This method assumes that upon the removal of any given parameter, the influence of the remaining parameters on the objective function remains unchanged. Based on OBD, OBS (hassibi1993optimal) and OBQ (frantar2022optimal) calculate the impact of each parameter weight on the objective function by solving the inverse Hessian matrix. Concurrently, they compute a compensation term applied to the remaining weights to offset the error introduced by each weight adjustment.
Although one-by-one weight quantization methods have achieved satisfactory performance on smaller models, the computational overhead becomes prohibitive when scaling to larger models. To accelerate quantization, GPTQ (frantar2022gptq) quantizes the weights column-by-column, and the rounding errors are compensated using second-order information. Specifically, this algorithm compensates for the quantization error induced by the quantized weights $Quant(\mathbf{W}_{i})$ by adjusting the subset R of full-precision weights $R$ with an update $\boldsymbol{\delta}_{R}$:
@@
where the Hessian matrix is $\mathbf{H}_{R}=2\mathbf{X}_{R}\mathbf{X}_{R}^{\top}$. Based on GPTQ, several works have been successively proposed. QuantEase (behdin2023quantease) utilizes the Coordinate Descent to compute more precise compensation for the unquantized weights. QQQ (zhang2024qqq) adopts the GPTQ for the transferred weights by OS+ (wei2023outlier).
@@ 5.2.3 Mixed-precision
As aforementioned, the presence of outliers is widely found in the activations and weights of large language models, which poses a significant challenge for quantization. Consequently, the motivation of numerous mixed-precision methods for LLMs is to represent a small number of outlier values in higher precision and other values in lower precision separately. Similarly, depending on the granularity of mixed precision, methods can be categorized into element-wise, channel-wise, token-wise and tensor-wise approaches, as described in Section 2.2.
@@
Element-wise. SpQR (dettmers2023spqr) was the first to demonstrate that outliers also exist in weights. It identifies and isolates these outlier weights based on their sensitivity, saving them as a highly sparse, higher-precision matrix. SqueezeLLM (kim2023squeezellm) adopts non-uniform quantization for non-salient weights, which achieves near-lossless performance. Similarly, CherryQ (cui2024cherry) defines heterogeneity to identify the critical cherry parameters. To explore extreme compression rate, PB-LLM (shang2023pb) is the first to binarize the non-salient weights in LLMs. Since PB-LLM still allocates high precision to 10%-30% of salient weights, BiLLM (huang2024billm) employs residual approximation for salient weights and group quantization for non-salient weights, reducing the quantization bit-width of LLM weights to 1.08 bits. GEAR (kang2024gear) extends the concept of mixed precision to the KV cache compression and utilizes low-rank matrices to approximate the quantization residuals.
@@
Channel-wise. LLM.int8() (dettmers2022gpt3) splits the weights and activations into two independent parts according to the outlier channels to minimize output quantization errors in activations, which effectively reduces the GPU memory usage during inference. OWQ (lee2024owq) proposes a sensitivity-aware mixed-precision scheme to identify the weak columns by Hessian metric. Furthermore, OWQ also provides weak column tuning (WCT) to enable accurate parameter-efficient fine-tuning for task-specific adaptation. RPTQ (yuan2023rptq) observes the varying ranges across channels in activations pose challenges for quantization. Therefore, RPTQ reorders the channels into different clusters with respective quantization. Atom (zhao2024atom) employs dynamic reorder for activations and static reorder for weights to remain aligned with the corresponding activation channels. Atom further quantizes the KV cache to 4-bit which significantly boosts serving throughput. Inspired by information theory, CQ (zhang2024kv) couples multiple key/value channels together and jointly quantize them.
Token-wise. Some KV cache quantization studies, such as KVQuant (hooper2024kvquant), IntactKV (liu2024intactkv) and SKVQ (duanmu2024skvq) discover token-wise outliers caused by special tokens (first token or low-semantic-value tokens) significantly influence the performance. So they store these token-wise outliers with higher precision in advance. KIVI (liu2024kivi) and WKVQuant (yue2024wkvquant) keep the most recent KV cache in full-precision and quantize the past KV cache. MiKV (yang2024no), Zipcache (he2024zipcache), and Snapkv (li2024snapkv) retain the important KV pairs in high precision based on distinct metrics. QAQ (dong2024qaqqualityadaptivequantization) dynamically allocates the adaptive bits for the different tokens.
Tensor-wise. LLM-MQ (li2023llm) assigns higher bit-widths to more sensitive layers based on first-order information and quantization error. CacheGen (liu2024cachegen) identifies LLM is more sensitive to losses in the KV cache values of the early layers than to losses in those of the deeper layers. It assigns higher-bit precision in sensitive early layers. The QuantMoE-Bench (li2024examining) investigates the weight bits among different blocks, experts, and linear layers, revealing that the varying numbers of weight bits are effective.
@@ 5.2.4 Combination
Although QAT is generally considered to offer the best accuracy, its high memory cost makes it difficult to apply to LLMs. Therefore, some methods consider introducing LoRA or other matrix decomposition methods as a trade-off between PTQ and QAT. Unlike PEFT discussed in Section 3.3, these methods aim to reduce quantization error using techniques like LoRA or SVD to achieve a quantized model closer to the full-precision model, rather than enhancing learning ability on fine-tuning datasets. Some works have used LoRA to achieve parameter-efficient QAT. LR-QAT (bondarenko2024low) computes $s\cdot\mathrm{clamp}(\textbf{W}_{q}+\textbf{A}^{i\times r}\textbf{B}^{r\times o})$ during the forward pass and does not update W during the backward pass, allowing a 7B LLM to be trained on a single consumer-grade GPU with 24GB of memory. This approach results in a quantization-friendly model after fine-tuning. LLM-QFA aims to produce models with various bit widths through a single supernet training, significantly reducing the resource overhead of this production method by leveraging the low resource cost of LoRA. INT2.1 (chai2023int2) utilizes LoRA to shift the optimization target from minimizing per-layer or per-block quantization error to minimizing the overall output error of the model. Through end-to-end fine-tuning, it reduces the distance between the output distribution and its corresponding original full-precision counterpart. Other works have reduced quantization errors through matrix decomposition. LQER (zhang2024lqer) applies SVD to quantization errors and uses an activation-induced scaling matrix to guide the singular value distribution toward the desired pattern. Delta-CoMe (ping2024delta) discovers that the singular values of delta weights exhibit a long-tailed distribution after applying SVD, and proposes a mixed-precision delta quantization method that uses high-bit representations for the singular vectors corresponding to these singular values. ZeroQuant-V2 (yao2023zeroquant) introduced an optimized low-rank compensation method that enhances model quality recovery by leveraging a low-rank matrix obtained through SVD of the quantization errors. LCQ (cai2024lcq) uses low-rank codebooks with a rank greater than one for quantization, addressing the issue of accuracy loss when using rank-one codebooks under high compression ratios.
@@
Model sparsification aims to remove unimportant weights to accelerate the model, while quantization further reduces the remaining weights using lower-bit representations. Therefore, the two methods can be effectively used in a complementary manner. SDQ (jeong2024sdq) first sparsifies the weights of LLMs based on the magnitude as much as possible until the quality of the LLM is significantly impacted (e.g., a 1% increase in perplexity). Then it utilizes a mixed-precision quantization method to deal with the outliers. However, this method does not take into account the coupling of the two approaches. Sparsification and quantization often conflict with each other. Sparsification tends to preserve parameters with large absolute values in LLMs (han2015deep; sun2023simple), while quantization prefers a smaller range of parameter values (wei2023outlier). As a result, the parameters preserved during sparsification may degrade the performance of quantization. JSQ (guo2024compressing) design a new sparsity metric to address this issue:
[…]
@@
$where\qquad\hat{\textbf{Y}}=\textbf{X}\cdot(\Theta(\textbf{W};i;j))^{\mathsf{T}},$
@@
In addition to combining quantization with other compression methods, different quantization techniques can also be integrated to achieve better results. A recent work (sharify2024combining) combines the SmoothQuant and GPTQ together. Actually, most of the equivalent transformation methods and the compensation quantization methods are orthogonal which can be merged for further exploration.
5.2.5 More LLM-Based Architectures
Besides traditional dense LLMs, the quantization methods tailored for multimodal large language models (MLLMs) and mixture-of-expert (MoE) models have also garnered widespread attention. Q-VLM (wang2024q) offers the first post-training quantization framework for MLLMs by mining cross-layer dependency to achieve satisfying trade-offs between discretization errors and the search cost. MQuant (yu2025mquant) proposes a static solution, utilizing seperate quantization parameters for visual and language modality. Furthermore, it relieves weight outliers arising from online Hadamard rotations. MBQ (li2024mbq) also considers the sensitivity between language and vision modality, which adjusts the reconstruction loss for the optimal channel-wise equalization factors. QuantMoE-Bench (li2024examining) explores the structure-aware mix-precision quantization schemes for MOE models, indicating different MoE structures require varying numbers of bits. MC-MOE (huang2024mc) converts the bit allocation problem into a Linear Programming (LP) problem and balances the importance between each expert.
@@ 5.2.6 More Quantization Forms
Integer quantization typically assigns a single scaling factor to an entire block and quantizes each element individually into an integer number. This reduces memory usage while also enabling the acceleration of fixed-point operations after quantizing both weights and activations. However, as higher precision is demanded for LLM quantization, formats that better match the original distribution of values have been proposed. Normal Float (dettmers20218; dettmers2024qlora), proposed alongside Quantile Quantization, is based on the assumption that the weight distribution follows a normal distribution. It is considered an information-theoretically optimal data type that ensures each quantization bin has an equal number of values assigned from the input tensor. However, Dotzel et al. (dotzel2024learningstudentsapplyingtdistributions) conducted a statistical analysis and found that the distributions of most LLM weights and activations follow a Student’s t-distribution. Based on this, they derived a new theoretically optimal format, Student Float (SF4). Floating-Point (FP) quantization offers better hardware support compared to NF/SF and is more flexible than integer quantization, allowing it to more effectively handle long-tail or bell-shaped distributions. Since FP can support flexible allocation of exponent and mantissa bits, several allocation schemes have been proposed. FPQ (liu2023llm) determines FP quantizers through a joint format and max value search combined with a pre-shifted exponent bias. FP8 quantization (kuzmin2022fp8) tests various allocation schemes by evaluating metrics like quantization error and proposes FP8 quantization simulation for learnable allocation and quantization.
@@
Transformer-VQ (lingle2023transformer) applies vector quantization (VQ) to the key vector sequence of Attention, reducing the complexity of Attention to linear. Most other VQ works focus on optimizing the codebooks $C_{m}\in\mathbb{R}^{2^{B}}$, and the discrete codes represented by one-hot $d$. AQLM (egiazarian2024extreme) learns additive quantization of weight matrices in an input-adaptive fashion and jointly optimizes codebook parameters across each transformer block. QuIP# (egiazarian2024extreme) uses vector quantization to exploit the spherical sub-Gaussian distribution inherent in incoherent weights by introducing a hardware-efficient codebook based on the highly symmetrical E8 lattice. GPTVQ (van2024gptvq) interleaves the quantization of one or more columns with updates to the remaining unquantized weights, using information from the Hessian of the per-layer output reconstruction MSE, and further compresses the codebooks by using integer quantization and SVD-based compression. PV-Tuning (malinovskii2024pv) notes that using straight-through estimators (STE) leads to suboptimal results and proposes an alternating iterative optimization strategy for scales, codebooks, zeros (continuous parameters), and assignments (discrete codes) during fine-tuning. QTIP (tseng2024qtip) uses a stateful decoder that separates the codebook size from the bitrate and effective dimension to achieve ultra-high-dimensional quantization.
Takeaways of subsection 5.2 For standard Post-Training Quantization (PTQ) of LLMs, equivalent transformation techniques such as shifting, scaling, and rotation can be employed to mitigate the impact of outliers. Quantization error can be further minimized through advanced compensation methods like GPTQ. For scenarios prioritizing high accuracy, mixed-precision quantization can be applied to recover performance loss. Conversely, if a high compression rate is the goal, combining low-rank approximation and sparsity-based methods can be effective. Furthermore, there are unique opportunities to explore emerging data formats, novel quantization functions, and cutting-edge model architectures, such as Multimodal Large Language Models (MLLMs) and Mixture of Expert (MOE) models.